OpenStudy (anonymous):

What is the slope-intercept form of the equation of a line that passes through (5, -4) and has a slope of 34? y=−34x−14 y=34x+14 y=34x−314 y=34x+314 Done

OpenStudy (texaschic101):

slope intercept form is y = mx + b m is your slope and b is your y intercept we already know the slope (m) is 34....and the points (5,-4)...x = 5 and y = -4. So basically, we just need to find the y intercept (b) now we use the formula y = mx + b .....now we sub out info in and solve for b -4 = 34(5) + b -4 = 170 + b -4 - 170 = b -174 = b so your equation is : y = 34x - 174 are you sure the problem is written correctly ??

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

but it does not match your answer choices

OpenStudy (anonymous):

can u help me on a different question my teacher helped me but she can only help me on one

OpenStudy (anonymous):

A company that packages cold cuts needs to transport the packaged meat to the supermarket. It will cost $1,200 to rent a truck, and it will cost an additional $100 for every 50 pounds transported. Which equation represents the transport cost for every 50 pounds of cold cuts transported? y = 1,200 − 100x y = 1,200 + 100x y = 1,200x + 100 y = 100x − 1,200

OpenStudy (texaschic101):

y = 1200 + 100x (y) transport cost = 1200 for rent of truck + 100x (x being however many 50 lbs of meat there is)